OpenStudy (anonymous):

The two-way table shows the number of men and women who voted for Candidate A or B. What percent of people who voted for Candidate A were women? Express your answer to the nearest whole percent. % Men Women Candidate A 1253 1921 Candidate B 1564 1723

OpenStudy (kamibug):

1253 + 1921 = 3174 1921 / 3174 = approx. 0.605 0.605 * 100 = ?